OpenStudy (anonymous):

Help please. I'll give a medal and become a fan! Find the foci of this ellipse equation. 4x^2+9y^2=36.

OpenStudy (anonymous):

lol no one likes these conic section questions but that are not that bad first divide everything by 36

OpenStudy (anonymous):

\[\frac{x^2}{9}+\frac{y^2}{4}=1\]

OpenStudy (anonymous):

which looks a lot like \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

you good from there? center is \((0,0)\) for example and since \(3^2=9\) the vertices are at \((-3,0)\) and \((3,0)\)

OpenStudy (anonymous):

oh you need the foci also that is ok because we have \(a=3\) and \(b=2\) so what is missing is \(c\) and \(c^2=a^2-b^2\)

OpenStudy (anonymous):

So to find the foci I do c^2=a^2-b^2 c^2=9^2-4^2 c^2=81-16 c^2=65 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[c^2=a^2-b^2\\ c^2=9-4\\ c^2=5\]

OpenStudy (anonymous):

your denominators are the \(a^2\) and \(b^2\) you don't square them again

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

i always get confused with that

OpenStudy (anonymous):

before we finish, is it clear that the center is \((0,0)\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok you need to know the center to find the foci you also need to know something else

OpenStudy (anonymous):



OpenStudy (anonymous):

is it the one on the left, or the one on the right?

OpenStudy (anonymous):

I believe it's the second one. I hope.

OpenStudy (anonymous):

yes, and you know that because in \[\frac{x^2}{9}+\frac{y^2}{4}=1\] the larger number is under the \(x\)

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

but you need that because now you know the foci are \(\sqrt5\) units to the RIGHT AND LEFT of the center, not UP and DOWN

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and now we are done right?

OpenStudy (anonymous):

i believe so.

OpenStudy (anonymous):

(+/- square root 5, 0)?

OpenStudy (anonymous):

Thanks @satellite73